# Reactive Diffusion Policy: Slow-Fast Visual-Tactile Policy Learning for Contact-Rich Manipulation

### v1

@@ Abstract
Figure 1: TactAR is a low-cost and versatile teleoperation system which can provide real-time tactile / force feedback via Augmented Reality (AR). Reactive Diffusion Policy (RDP) is a slow-fast imitation learning algorithm that can model complex behaviors with a slow policy network and achieve closed-loop control based on tactile / force feedback with a fast policy network.
@@ I Introduction
  • A teleoperation system called TactAR which can provide fine-grained tactile feedback in real time through Augmented Reality (AR).
@@
Figure 2: Overview of TactAR teleoperation system. It can provide real-time tactile / force feedback via Augmented Reality (AR). The tactile feedback is represented as the 3D deformation field, which is a universal representation applicable to multiple different tactile / force sensors. The 3D deformation field is rendered and ”attached” to the robot end-effector in AR, which makes the user perceive the rich contact information in 3D space. TactAR also support real-time streaming for multiple RGB cameras and optical tactile sensors. Please see the video in the supplementary file for more details.
@@ II-B Robot Data Collection System
Some research works use VR headset to get tactile feedback. Comp-ACT [34] and DIPCOM [1] use VR controller to adjust the stiffness of the low-level robot controller during teleoperation, but it cannot provide active force feedback and relies on robot arms with built-in torque sensors. Zhou et al. [73] and Bunny-VisionPro [14] use vibration to provide tactile / force feedback, but they can only provide very coarse information about contact. Our method combines the advantages of low-cost VR controller and tactile sensing, getting tactile feedback via Augmented Reality, while preserving the accuracy needed for precise contact-rich tasks. In addition, our teleoperation systems are flexible and easy to deploy on different tactile sensors and robot platforms.
@@ II-C Visual-tactile Manipulation
Tactile sensing is crucial in robotic manipulation, compensating for visual occlusions and providing force/torque feedback in complex tasks. While tactile perception has a long history [66, 68, 25, 2], many approaches rely on task-specific modeling [45], hand-designed primitives [66], or hand-crafted rewards [68], limiting their generalizability. For instance, TAVI [25] uses uSkin [57] sensors and online RL driven by visual rewards but is constrained to tasks with minimal wrist movement due to safety concerns. Similarly, RoboPack [2] employs SoftBubble [37] tactile sensors and MPC but struggles with slow planning speeds in contact-rich tasks.
Recent advances in Visual Imitation Learning (e.g., [11, 71, 69]) have shown promise in learning complex tasks end-to-end. Several studies [42, 29, 8, 65] have attempted to integrate tactile sensing with imitation learning, but most focus solely on normal force readings, limiting their applicability to versatile tasks. For example, HATO [42] uses sparse tactile sensors for bimanual manipulation but lacks the resolution and shear force for more precise, contact-rich tasks. 3D-ViTac [29] adds tactile arrays to the ALOHA [71] gripper but only captures normal pressure, restricting its use to tasks with visual occlusion or in-hand force control.
Few works leverage shear force/torque data for imitation learning. MimicTouch [65] uses a GelSight [67]-based handheld device for offline policy training but omits visual input due to the embodiment gap between human hands and robot grippers, limiting its task versatility. Similarly, [8] uses GelSlim [53]’s shear force field for policy transfer across robot embodiments but also excludes visual input, restricting their approach to simple tasks. In contrast, our method combines normal force, shear force, and visual RGB inputs into a unified visual-tactile policy, enabling deployment across a broader range of tasks. By integrating both tactile and visual modalities, our approach overcomes the limitations of prior works and achieves greater versatility in robotic manipulation.
@@ III Teleoperation System: TactAR
Previous works [34, 9, 14, 28] have demonstrated the importance of incorporating tactile/haptic feedback during teleoperation. For optical tactile sensors, one straightforward way to provide tactile feedback is to directly visualize the raw tactile images on the screen. However, we have found that directly presenting 2D tactile images does not enable operators to intuitively perceive force directions and torques in 3D space. Thus, we choose to render the 3D deformation field of the marker motion in Augmented Reality (AR) and attach it to the robot end-effector in virtual space. In this way, the operator can directly perceive the rich contact information of the tactile sensor (e.g., the magnitude and direction of normal force, shear force and torque).
The marker array on the tactile optical sensors [67, 56, 50](see Fig. 3) can track the deformation of gel surface, which provides rich information about contact (i.e., normal force , tangential force and torques). However, directly calculating the force/torque information from the 2D marker flow field requires complex calibration process with expensive force/torque sensors [70, 21]. In order to enhance the accessibility of our system, we choose to directly visualize the 3D marker deformation field in AR. We denote the original image in frame $t$ captured by the optical tactile sensor as $I_{t}$. We first extract the normalized 2D location of the marker arrays $D_{t}$ with OpenCV [46] from the image $I_{t}$. Then we use a score-based tracking algorithm [23] to calculate the 2D optical flow from the undeformed frame $D_{0}$ to the current frame $D_{t}$ as Eq 1:
$F_{t}=[\mathbf{d}_{x},\mathbf{d}_{y}]=Flow(D_{0},D_{t})$ (1)
We can get 3D deformation field $V_{t}=[\mathbf{d}_{x},\mathbf{d}_{y},\mathbf{o_{z}}]$ used in AR, here $\mathbf{o_{z}}$ denotes the z-axis offsets for 3D visualization. For force sensors, we can directly get $V_{t}=[\mathbf{f}_{x},\mathbf{f}_{y},\mathbf{f_{z}}]$ for 3D visualization.
@@ III-B Real-time Tactile / Force Feedback Rendering in AR
In our TactAR system, we use the full-color pass-through mode of the Meta Quest3 VR headset to create an immersive AR world which is implemented by Unity. The Meta Quest 3 VR headset has a native refresh rate of 90Hz, and it has a bulit-in SLAM algorithm for pose estimation of the headset and controllers. Before teleoperation, we first align the AR coordinate system in Quest3 with the real-world robot coordinate system by a simple camera calibration process described in Fig. 4.
As shown in Fig. 2, our TactAR system consists of a workstation node, RGB camera nodes, tactile camera nodes, robot controller node and the VR headset node. We use ROS2 to perform time synchronization between different sensor topics. The robot controller node will publish the TCP pose (Tool Center Point) and the tactile camera nodes will publish the 3D deformation field $V_{t}$ defined in Sec. III-A. We transform the 3D deformation field $V_{t}$ based on robot TCP pose and render it in AR space to allow the 3D deformation field to ”attach” to the TCP position at the robot end effector. In this way, the user can get rich contact information including the normal force, tangential force, and the torsional torque in 3D space. Please see Appendix -B and the video in the supplementary files for more details.
Our system can achieve low-latency feedback for tactile observation. Typically, the latency of the marker flow tracking algorithm is about 10ms. The latency for the force sensor is less than 1ms. The rendering latency in Quest 3 is about 10ms. And the network latency is about 1-6ms depending on the network condition. Optionally, our TactAR system also supports real-time streaming of multi-view RGB cameras (see in Fig. 2) and tactile cameras for more immersive teleoperation experience.
@@ III-C Versatility and Accessibility
  • Cross-sensor. Our TactAR system uses the 3D deformation field as the unified representation for tactile / force feedbacks, which can be applied on many different tactile / force sensors. In this paper, we the following sensors for experiments: (1) GelSight Mini [22] (Robotics Package) optical tactile sensor. (2) Our improved MCTac [50] optical tactile sensor. (3) Built-in joint torque sensors in Flexiv Rizon [20] robot arm.
    Our TactAR system is highly modularized and can even support AR visualization with different tactile / force sensors at the same time (see Fig. 1). Theoretically, our teleoperation system is also compatible with electrically tactile sensors like AnySkin [6], whose data types can typically be represented as a sparse 3D force field. This format perfectly aligns with our 3D deformation field representation.
@@
  • Low cost. Our TactAR system is built with low-cost hardwares. Our experiments use two different optical tactile sensors: (1) GelSight Mini Robotics Package [22] which is a commercialized product costs \$$\$549$. (2) Our customized MCTac [50] costs around \$$\$50$ in lab fabrication, and there is large potential for lower cost in industrial manufacturing in the future. Compared to force/torque sensors like ATI mini45 [3] which costs about \$$\$3000$, the optical tactile sensors offer a significant cost advantage. The Meta Quest3 VR heaset used for teleoperation and AR feedbacks costs \$$\$499$.
@@ IV-A Tactile / Force Representation
For an optical tactile sensor with $n$ marker dots, we can calculate the 2D deformation field matrix $F$ with shape $(n,2)$ according to Eq. 1 for each frame. Previous works [8, 47, 10] has proven that low-dimensional tactile representation based on deformation field/force field are easier to learn for contact-rich tasks compared to the raw high-dimensional tactile image. Besides, the deformation field is more robust and is less affected by noise such as differences in lighting and texture when the gel on the sensor surface is replaced or slightly damaged during evaluation.
According to Zhang et al. [70], the deformation field of optical tactile sensors can be decomposed into several independent components that are highly explicable (e.g., normal force, torsional force etc). Inspired by this, we use principal component decomposition (PCA) to reduce the dimension for the deformation field $F$. When encountering a new frame $F_{t^{\prime}}$, we can get the $d$-dim tactile PCA feature with the PCA projection matrix. In practice, we find $d=15$ is enough for reconstruction without losing much details. Please see the Appendix for more details of the PCA dataset (Appendix -D) and PCA reconstruction results (Appendix -C). We also visualize the first few main components of PCA and surprisingly find that these components are highly explicable (see Appendix -C) and represent the force/torque of different directions. The use of PCA feature makes the model more robust to tracking errors and noise of marker deformation field. The interpretability also allows the policy network to utilize the tactile representation more easily.
@@ IV-B Slow-Fast Policy Learning
Previous works [11, 72] have demonstrated that predicting action sequences or action chunks [38] effectively preserves temporal action consistency and handles non-Markovian or idle actions, which achieve superior performance in policy learning. However, when executing the action chunk, such approaches can be viewed as an open-loop policy, preventing it from achieving real-time feedback from high-frequency signals such as tactility. VISK [47] utilizes temporal ensembling to mitigate this issue. As shown in Fig. 5 (right), temporal ensembling finds a balance between closed-loop control and sequence consistency by aggregating the predictions of multiple iterations for the same timestep. A significant drawback of this solution is that it diminishes the policy’s ability to model multi-modal distributions and non-Markovian actions, making it prone to issues such as getting stuck. Moreover, we find that the policy performance is quite sensitive to the smoothing coefficient of temporal ensembling, which consequently reduces the policy’s applicability.
@@ IV-B2 Slow Policy
TABLE I: Inference Time of Different Modules
Diffusion Policy Slow Policy (LDP) Fast Policy (AE)
@@ IV-B3 Implementing Suggestions for Slow-Fast Policy
  • Relative trajectory. We use relative end-effector (EE) trajectory for action representation, proprioception, and intergripper proprioception, which has been proven to be effective even in complex tasks by UMI [12]. Specifically, instead of directly calculating the delta action between consecutive frames (may lead to large compounding errors), we convert an absolute pose trajectory to a relative one by calculating the relative transformation with respect to a base frame. In our setting, the base frame is the last observation frame in an action chunk.
@@ V Experiments
  • Q6: How does tactile / force feedback in TactAR contribute to the data quality in human demonstration?
@@ V-A4 Evaluation Protocols
We use similar initial states across all methods for both the robots and the objects, by manually aligning the scene with the pre-defined images. There are three test-time variations for Peeling and Wiping tasks: (a) No perturbation. The object is fixed with a random 6D pose in the air. (b) Perturbation before contact. The human evaluator will move the object right before the tool makes contact. (c) Perturbation after contact. The human evaluator will move the object after the tool makes contact to break the contact state. There are two test-time variations for Bimanual Lifting task: (a) soft paper cup. (b) hard paper cup. We run 10 trials for each test-time variation. Please see Appendix -E for the detailed definition of score and success rate calculation for each task.
@@ V-A5 Implementation Details
The Diffusion Policy and our slow policy (LDP) predict open-loop 12 FPS action sequences for each action chunk. The fast policy (AE) takes tactile / force observations at 24 FPS and output action predictions at 24 FPS. The final actions are interpolated and sent to robots with a higher frequency ($>$500Hz). Note that we use 24 FPS because we are constrained by the frame rate limitation of GelSight [22], which is 25 FPS. Our RDP algorithm can also be applied to higher frequency tactile / force signals in theory. Please see Appendix -F for more details on data collection and the inference process.
@@ V-B Results
Simply adding tactile signals into observation may NOT improve performance for complex contact-rich tasks (Q1). As shown in Tab. II , compared to low-dim tactile embedding (DP w. tactile emb.), simply using raw tactile images of GelSight (DP w. tactile img.) can significantly degrade the model performance of Diffusion Policy ($0.39\to 0.03$). A possible explanation for this phenomenon is that high-dimensional image inputs are more susceptible to being out of distribution (OOD) when the dataset size is small.
We also found that Diffusion Policy (DP) incorporates tactile embedding (DP w. tactile emb.) performs similarly compared to DP with purely visual input in the three tasks (see Tab II, Tab III and Tab IV). However, despite similar performance, these two DP baselines exhibit different failure modes. We observe that DP with pure visual input frequently predicts inaccurate trajectories and results in large contact forces (e.g., failure case 2 in Fig. 8 (b)), necessitating the human evaluators to move their hands to prevent sensor damage. In contrast, DP with tactile embedding rarely brings large contact forces. It may get stuck when making contact with the object (e.g., failure case 2 in Fig. 8 (a) & failure case in Fig. 8 (c)). This may be because it learns some reactive behavior from the data (e.g., slightly move up when making contact; move down when loose contact). However, since it executes action chunks in an open-loop manner, it lacks the capability for finer adjustments. Consequently, it might quickly transition to a different contact state and repeatedly switch between various contact states. In conclusion, we need a better way to incorporate tactile signals into the policy.
@@
We have performed some case studies and visualization to analyze how RDP works during these contact-rich tasks in Fig. 9. We have observed that RDP indeed learns reactive behaviors similar to those of humans. For instance, in Case Study 1, when a peeler suddenly comes into contact with a cucumber, the contact force increases abruptly. The fast policy then applies an upward adjustment to the action chunk predicted by the slow policy to reduce the contact force. In Case Study 2, as the peeler approaches the end of the cucumber, the fast policy predicts a downward corrective action, allowing the peeler to closely follow the cucumber’s surface and remove more skin. In Case Study 3, when two handlers attempt to clamp the paper cup, the fast policy swiftly predicts an outward corrective action once contact is made, preventing the cup from being squeezed. The magnitude of these reactive behavior adjustments is at the sub-millimeter level, which might appear small but significantly impacts the overall performance of these contact-rich tasks. These minor reactive actions are hard to learn accurately for a policy with open-loop action chunking, but much easier to learn for our fast policy with high-frequency tactile input.
@@
RDP are applicable to different tactile / force sensors (Q3). As shown in Tab. II, the performance of RDP with the two optical tactile sensors (GelSight Mini [22] and MCTac [50]) are very close (0.9 vs 0.88) on Peeling task. Furthermore, we surprisingly find that RDP can perform well using different tactile sensors (MCTac on the left gripper and GelSight on the right gripper) at the same time for the Bimanual Lifting task (see Tab. IV). Note that GelSight Mini and MCTac have different LED colors (RGB vs white), different gel material (hard vs soft), different marker arrays ($7*9$ vs $5*7$), different resolutions (8MP vs 2MP) and different frame rates (25 FPS and 30 FPS). This validates the effectiveness of our tactile representation and the versatility of the RDP algorithm.
@@
RDP can react immediately to external disturbances (Q4). We find that the slow-fast design of RDP significantly improved the model’s response speed to external disturbances. For example, in Tab. II, RDP (GelSight) achieves a score of 0.79 on the hardest setting (Perturbation after Contact) in Peeling task, while the DP baseline with tactile embedding only achieves a score of 0.15. In the Wiping task (see Tab. III), we find that RDP also achieves a higher score compared to DP baselines under human perturbations. Fig. 8 has shown some typical failure cases for DP baselines under human perturbations (e.g., failure case 1 in Peeling (a) task and Wiping (b) task). When the robot loses contact with perturbations (e.g., moving down), the DP baselines will execute the remaining trajectory in the action chunk in an open-loop manner, resulting in broken cucumber peels and residual handwriting on the vase that were not completely wiped off. In contrast, the RDP algorithm can immediately change the predicted trajectory with the fast policy in a closed-loop manner, which leads to better results.
Slow-fast hierarchy, relative trajectory and latency matching are essential for RDP performance (Q5). As shown in Fig. 5 (right), there are two ways to increase the closed-loop control frequency without our slow-fast hierarchy: (1) reducing action chunk size. (2) using temporal ensemble. However, experiments in Tab. V have proved that these two options both have significant side effects. We can see from Tab. V that when the action chunk size is reduced from 8 to 2, the DP baseline tends to get stuck before grasping (failure case 4 in Fig. 8 (b)), which makes the grasp success rate drop from $100\%$ to $20\%$. Policy with small chunk size is very sensitive to non-markovion behaviors (e.g., pauses in the air) commonly found in human demonstration data, so we can not simply reduce chunk size.
@@
Figure 11: Data Collection Quality Comparison. Data collected with TactAR has higher stability of contact forces compared to traditional VR teleoperation.
Tactile / force feedback in TactAR improves data quality in contact-rich tasks by improving the stability of contact forces (Q6). We use two different teleoperation settings to collect data (10 demos) for Peeling task with the same volunteer: (1) VR (TactAR without tactile / force feedback) (2) TactAR with tactile / force feedback, then we calculate the Rolling Standard Deviation of the recorded force curve with a window size of 10 steps, and the results are shown in Fig. 11. We can observe that using TactAR to collect data helps avoid large rolling standard deviation, indicating reduced temporal fluctuations. Data with higher stability of contact forces will help the model to discover useful patterns more easily. From an alternative perspective, the excellent performance of our RDP model also indirectly validates the effectiveness of the data collected by TactAR.

### v2

@@ Abstract
Figure 1: TactAR is a low-cost and versatile teleoperation system which can provide real-time tactile / force feedback via Augmented Reality (AR). Reactive Diffusion Policy (RDP) is a slow-fast imitation learning algorithm that can model complex action trajectories with a slow policy network and achieve closed-loop control based on high-frequency tactile / force feedback with a fast policy network.
@@ I Introduction
  • A teleoperation system called TactAR which can provide fine-grained tactile / force feedback in real time through Augmented Reality (AR).
@@
Figure 2: Overview of TactAR teleoperation system. It can provide real-time tactile / force feedback via Augmented Reality (AR). The tactile feedback is represented as the 3D deformation field, which is a universal representation applicable to multiple different tactile / force sensors. The 3D deformation field is rendered and “attached” to the robot end-effector in AR, which makes the user perceive the rich contact information in 3D space. TactAR also support real-time streaming for multiple RGB cameras and optical tactile sensors. Please see the video in the supplementary file for more details.
@@ II-B Robot Data Collection System
Some research works use vibration in VR headset [32, 1, 72, 13] to get tactile feedback, but these systems can only provide very coarse information about contact. Our method combines the advantages of low-cost VR controller and tactile sensing, getting tactile feedback via Augmented Reality, while preserving the accuracy needed for precise contact-rich tasks. In addition, our teleoperation systems are flexible and easy to deploy on different tactile sensors and robot platforms.
@@ II-C Visual-tactile Manipulation
Tactile sensing is crucial in robotic manipulation, compensating for visual occlusions and providing force/torque feedback in complex tasks. While manipulation with tactile sensing has a long history [65, 67, 23, 2], many approaches rely on task-specific modeling [43, 2], hand-designed primitives [65, 63], or hand-crafted rewards [67], limiting their generalizability on different tasks.
Recent advances in visual Imitation Learning (e.g., [10, 70, 68]) have shown promise in learning complex tasks in an end-to-end way. Several studies [40, 27, 8, 64] have attempted to integrate tactile sensing with imitation learning, but most focus solely on normal force readings, limiting their applicability to versatile tasks. Only a few works leverage shear force/torque data for imitation learning. MimicTouch [64] uses a GelSight [66]-based handheld device for offline policy training but omits visual input due to the embodiment gap between human hands and robot grippers, limiting its task versatility. Similarly, Bogert et al. [8] uses GelSlim [51]’s shear force field for policy transfer across robot embodiments but also excludes visual input, restricting their approach to simple tasks.
In contrast, our method combines normal force, shear force, and visual RGB inputs into a unified visual-tactile policy, enabling deployment across a broader range of tasks. By integrating both tactile and visual modalities, our approach overcomes the limitations of prior works and achieves greater versatility in robotic manipulation.
@@ III Teleoperation System: TactAR
The gel surface’s marker array (Fig. 3) captures rich contact information, but deriving forces from 2D optical flow requires complex calibration with expensive sensors. To improve accessibility, we instead visualize the 3D deformation field. From tactile images $I_{t}$, we extract normalized marker positions $D_{t}$ using OpenCV [44]. We use a score-based tracking algorithm [21] to calculate 2D optical flow between the initial frame $D_{0}$ and the current frame $D_{t}$:
$F_{t}=[\mathbf{d}_{x},\mathbf{d}_{y}]=Flow(D_{0},D_{t})$ (1)
The 3D deformation field $V_{t}=[\mathbf{d}_{x},\mathbf{d}_{y},\mathbf{o}_{z}]$ (with z-offset $\mathbf{o_{z}}$) is then rendered in AR. For force sensors, we directly visualize $V_{t}=[\mathbf{f}_{x},\mathbf{f}_{y},\mathbf{f}_{z}]$.
@@ III-B Real-time Tactile / Force Feedback Rendering in AR
Our TactAR system employs Meta Quest3’s color pass-through mode through Unity to create an AR environment. The native SLAM algorithm in Quest3 tracks headset/controller poses, with initial coordinate alignment between AR and robot spaces via a simple calibration process (Fig. 4). The ROS2-based architecture (Fig. 2) synchronizes data streams from tactile/force sensors, robot controllers, and RGB cameras. Tactile/force information (3D deformation field $V_{t}$) is transformed using real-time robot TCP poses and rendered in AR space. Optionally, our TactAR system also supports real-time streaming of multi-view RGB cameras (see in Fig. 2) and tactile cameras for more immersive teleoperation experience. Please see Appendix -B and our website for the system latency and more details of TactAR.
@@ III-C Versatility and Accessibility
  • Cross-sensor. Our TactAR system uses the 3D deformation field as the unified representation for tactile / force feedbacks, which can be applied on many different tactile / force sensors. In this paper, we the following sensors for experiments: (1) GelSight Mini [20] (Robotics Package) optical tactile sensor. (2) Our improved MCTac [48] optical tactile sensor. (3) Built-in joint torque sensors in Flexiv Rizon [19] robot arm. Our TactAR system is highly modularized and can even support AR visualization with different tactile / force sensors at the same time (see Fig. 1). Theoretically, our teleoperation system is also compatible with electrically tactile sensors like AnySkin [6], whose data types can typically be represented as a sparse 3D force field.
@@
  • Low cost. Our TactAR system is built with low-cost hardwares. The Meta Quest3 VR heaset used for teleoperation and AR feedbacks costs \$$\$499$. Our experiments use two different optical tactile sensors: (1) GelSight Mini Robotics Package [20] which is a commercialized product costs \$$\$549$. (2) Our customized MCTac [48] costs around \$$\$50$ in lab fabrication, and there is large potential for lower cost in industrial manufacturing in the future. Compared to force/torque sensors like ATI mini45 [3] which costs about \$$\$3000$, the optical tactile sensors offer a significant cost advantage.
@@ IV-A Tactile / Force Representation
For optical tactile sensors (e.g. Gelsight Mini [20]), we use low-dimensional representation generated by Principal Component Analysis (PCA) on the marker deformation field $\mathbf{F}$ in Eq. 1. The use of PCA feature makes the model more robust to tracking errors and noise of marker deformation field. In addition, PCA feature are more robust to texture and lighting changes due to damage or gel replacements. Please see Appendix -C and Appendix -D for more details of the tactile representation and the tactile dataset for PCA.
@@ IV-B Slow-Fast Policy Learning
Previous works [10, 71] have demonstrated that predicting action sequences or action chunks [36] effectively preserves temporal action consistency and handles non-Markovian or idle actions, which achieve superior performance in policy learning. However, when executing the action chunk, such approaches can be viewed as an open-loop policy, preventing it from achieving real-time feedback from high-frequency signals such as tactility. VISK [45] utilizes temporal ensembling to mitigate this issue. As shown in Fig. 5, temporal ensembling finds a balance between closed-loop control and sequence consistency by aggregating the predictions of multiple iterations for the same timestep. A significant drawback of this solution is that it diminishes the policy’s ability to model multi-modal distributions and non-Markovian actions, making it prone to issues such as getting stuck. Moreover, we find that the policy performance is quite sensitive to the smoothing coefficient of temporal ensembling, which consequently reduces the policy’s applicability.
@@ IV-B2 Slow Policy
TABLE I: Inference Time of Different Modules on RTX 4090
Diffusion Policy Slow Policy (LDP) Fast Policy (AT)
@@ IV-B3 Implementing Suggestions for Slow-Fast Policy
  • Relative trajectory. We use relative end-effector (EE) trajectory for action representation, which has been proven to be effective even in complex tasks by UMI [11]. Specifically, instead of directly calculating the delta action between consecutive frames (may lead to large compounding errors), we convert an absolute pose trajectory to a relative one by calculating the relative transformation with respect to a base frame. In our setting, the base frame is the last observation frame in an action chunk.
@@ V Experiments
  • Q6: How does tactile / force feedback in TactAR contribute to the data quality in teleoperation?
  • Q7: How does data quality in teleoperation influence policy performance?
@@ V-A4 Evaluation Protocols
We use similar initial states across all methods for both the robots and the objects, by manually aligning the scene with the pre-defined images. There are three test-time variations for Peeling and Wiping tasks: (a) No perturbation. The object is fixed with a random 6D pose in the air. (b) Perturbation before contact. The human evaluator will move the object right before the tool makes contact. (c) Perturbation after contact. The human evaluator will move the object after the tool makes contact to break the contact state. There are two test-time variations for Bimanual Lifting task: (a) soft paper cup. (b) hard paper cup. We run 10 trials for each test-time variation.
For Peeling task, we calculate the score based on the proportion of the peeled cucumber skin to the total length of the cucumber, normalized by the average score of the demonstration data. For Wiping task, we calculate the score based on the size of the remaining handwriting compared to the demonstration data. If the residue reaches the human demonstration level, the score is 1; If there is minor residue (less than one third of the handwriting length), the score is 0.5; If significant residue remains, the score is 0. For Bimanual Lifting task, if the paper cup is lifted into the air following the designated trajectory without significant compression, the score will be 1; If the paper cup is partially compressed in the air, the score will be 0.5; If the cup is not lifted up, or dropped in the air, the score will be 0. Please see Appendix -E for more details of the evaluation protocal.
@@ V-A5 Implementation Details
The Diffusion Policy and our slow policy (LDP) predict open-loop 12 FPS action sequences for each action chunk. The low-frequency observation of LDP includes both visual inputs and tactile / force inputs. The fast policy (AT) takes tactile / force observations at 24 FPS and update new action predictions at 24 FPS. Note that we use 24 FPS because we are constrained by the frame rate limitation of GelSight [20], which is 25 FPS. Our RDP algorithm can also be applied to higher frequency tactile / force signals in theory. Please see Appendix -D, -F and -H for more details on data collection, the inference process and the hyperparameters.
@@ V-B Results
Simply adding tactile signals into observation may NOT improve performance for complex contact-rich tasks (Q1). We have compared the performance of Diffusion Policy using raw tactile images (DP w. tactile img.) v.s. low-dim tactile embedding (DP w. tactile emb.) in Tab. II. Although the performance of both methods is similar, low-dimensional tactile embedding demonstrates greater robustness to texture changes resulting from gel damage or gel replacements during the evaluation process. Therefore, we use tactile embedding in most of our experiments.
We also found that Diffusion Policy (DP) incorporates tactile embedding (DP w. tactile emb.) performs similarly compared to DP with purely visual inputs in the three tasks (see Tab II, Tab III and Tab IV). However, despite similar performance, these two DP baselines exhibit different failure modes. We observe that DP with pure visual input frequently predicts inaccurate trajectories and results in large contact forces (e.g., failure case 2 in Fig. 8 (b)), necessitating the human evaluators to move their hands to prevent sensor damage. In contrast, DP with tactile embedding rarely brings large contact forces. It may get stuck when making contact with the object (e.g., failure case 2 in Fig. 8 (a) & failure case in Fig. 8 (c)). This may be because it learns some reactive behavior from the data (e.g., slightly move up when making contact; move down when loose contact). However, since it executes action chunks in an open-loop manner, it lacks the capability for finer adjustments. Consequently, it might quickly transition to a different contact state and repeatedly switch between various contact states. In conclusion, we need a better way to incorporate tactile signals into the policy.
@@
We have performed some case studies and visualization to analyze how RDP works during these contact-rich tasks in Fig. 9. We have observed that RDP indeed learns reactive behaviors similar to those of humans. For instance, in Case Study 1, when a peeler suddenly comes into contact with a cucumber, the contact force increases abruptly. The fast policy then applies an upward adjustment to the action chunk predicted by the slow policy to reduce the contact force. In Case Study 2, as the peeler approaches the end of the cucumber, the fast policy predicts a downward corrective action, allowing the peeler to closely follow the cucumber’s surface and remove more skin. In Case Study 3, when two handlers attempt to clamp the paper cup, the fast policy swiftly predicts an outward corrective action once contact is made, preventing the cup from being squeezed. The magnitude of these reactive behavior adjustments is at the sub-millimeter level, which might appear small but significantly impacts the overall performance of these contact-rich tasks. These minor reactive actions are hard to learn accurately for a policy with open-loop action chunking, but much easier to learn for our fast policy with high-frequency tactile / force control.
@@
RDP are applicable to different tactile / force sensors (Q3). As shown in Tab. II, the performance of RDP with the two optical tactile sensors (GelSight Mini [20] and MCTac [48]) are very close (0.9 vs 0.88) on Peeling task. Furthermore, we surprisingly find that RDP can perform well using different tactile sensors (MCTac on the left gripper and GelSight on the right gripper) at the same time for the Bimanual Lifting task (see Tab. IV). Note that GelSight Mini and MCTac have different LED colors (RGB vs white), different gel material (hard vs soft), different marker arrays ($7\times 9$ vs $5\times 7$), different resolutions (8MP vs 2MP) and different frame rates (25 FPS and 30 FPS). This validates the effectiveness of our tactile representation and the versatility of the RDP algorithm.
@@
RDP can react immediately to external disturbances (Q4). We find that the slow-fast design of RDP significantly improved the model’s response speed to external disturbances. For example, in Tab. II, RDP (GelSight) achieves a score of 0.8 on the hardest setting (Perturbation after Contact) in Peeling task, while the DP baseline with tactile embedding only achieves a score of 0.15. In the Wiping task (see Tab. III), we find that RDP also achieves a higher score compared to DP baselines under human perturbations. Fig. 8 has shown some typical failure cases for DP baselines under human perturbations (e.g., failure case 1 in Peeling (a) task and Wiping (b) task). When the robot loses contact with perturbations (e.g., moving down), the DP baselines will execute the remaining trajectory in the action chunk in an open-loop manner, resulting in broken cucumber peels and residual handwriting on the vase that were not completely wiped off. In contrast, the RDP algorithm can immediately change the predicted trajectory with the fast policy in a closed-loop manner, which leads to better results.
Slow-fast hierarchy, relative trajectory and latency matching are essential for RDP performance (Q5). As shown in Fig. 5, there are two ways to increase the closed-loop control frequency without our slow-fast hierarchy: (1) reducing action chunk size. (2) using temporal ensemble. However, experiments in Tab. V have proved that these two options both have significant side effects. We can see from Tab. V that when the action chunk size is reduced from 8 to 2, the DP baseline tends to get stuck before grasping (failure case 4 in Fig. 8 (b)), which makes the grasp success rate drop from $100\%$ to $20\%$. Policy with small chunk size is very sensitive to non-Markovion behaviors (e.g., pauses in the air) commonly found in human demonstration data, so we can not simply reduce chunk size.
@@
Tactile / force feedback in TactAR improves data quality in contact-rich tasks by improving the stability of contact forces (Q6). We have conducted a user study on how tactile / force feedback in TactAR helps the data collection process. We invited 10 users with different levels of experience in VR teleoperation or Imitation Learning (IL). Please see Appendix -G for more details of the user study. We perform $10\times 10\times 2=200$ trials in total for quantitative analysis. As shown in Fig. 11, most of the users ($\geq 70\%$) found that tactile / force AR feedback is very helpful in data collection, regardless of whether they were acting as the teleoperator or the person holding the cucumber. In Fig. 12, we can see that tactile / force feedback in TactAR can greatly improve the data quality from both the normalized peeling length ($0.72\rightarrow 0.91$) and the ratio of stable contact force ($0.58\rightarrow 0.87$).
Figure 11: User study results among 10 users on teleoperation w./w.o. tactile / force feedback in Peeling task.
Figure 12: Teleoperation data quality of 10 users on Peeling task (no perturb.) w./w.o. tactile / force feedback.
Figure 13: The stability of contact forces with different teleoperation systems in Peeling task. Data collected with TactAR has higher stability of contact forces compared to traditional VR teleoperation.
For a more detailed quantitative analysis of contact forces, we collect data (10 demos) for Peeling task with the same user by VR teleoperation and TactAR respectively, then we calculate the Rolling Standard Deviation of the recorded force curve with a window size of 10 steps, and the results are shown in Fig. 13. We can observe that using TactAR to collect data helps avoid a large rolling standard deviation, indicating reduced temporal fluctuations and more stable contact forces.
Figure 14: How data quality influences policy performance of RDP (force) in Peeling task.
Data with higher stability of contact forces will help the model to discover useful patterns more easily (Q7). We have collected the same number of demonstrations (60) for Peeling task with traditional VR teleoperation without tactile / force feedback and trained RDP (force) with these data. The results in Fig. 14 show that data quality has a large influence on policy performance (the score decreases by more than $30\%$). We observe that policies trained with low-quality data exhibited more unstable performance, such as more unstable force during peeling and a higher likelihood of breaking halfway through. A possible explanation is that the Fast Policy in RDP is designed to identify associations between tactile / force signals and trajectories from the data and learn reactive behavior. When contact forces in the data are highly unstable, the Fast Policy struggles to identify reasonable associations, which reduces performance.